A long clothesline is stretched horizontally between two trees. While shaking this clothesline up and down near one end at the rate of 4 cycles per second, the wavelength is 1m. What is the speed of these waves? If the distant end is 10m apart how long does a wave pulse take to return to you

1
Expert's answer
2022-01-07T09:29:24-0500

1.

"v=\lambda \nu=1\cdot 4=4~\frac ms."

2.

"t=\frac{2s}v=\frac{2\cdot 10}4=5~s."